OpenStudy (anonymous):

Find the exact area of a circle having the given circumference. 4(square root of three times pie) a= A. 2times the square root of three times pie B. 4 times the square root of three times pie C. 12 times pie

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

i got 24 as my final answer

OpenStudy (mathstudent55):

Your answer is not correct and is not in the choices. Divide the circumference by \(2\pi\) to find the radius, r. Then do \(\pi r^2 \) to find the area.

OpenStudy (nurali):

circumference is 2πr and here 2πr=4√3π so r=2√3π/π A=πr^2 A=π×(2√3π/π)^2 A=12π

OpenStudy (anonymous):

thanks guys

OpenStudy (nurali):

Anytime.